OpenStudy (anonymous):

Suppose a certain religious university has a ratio of approximately 90 males to 110 females enrolled as students. If the current enrollment at the university is 16,000 and every male on campus asks a female student to a dance, how many women will be without dates to the dance?

OpenStudy (anonymous):

Using guess and check, multiply 90 and 110 by the same number so that when you add the products up, it equals 16000. Once you find the total amount of boys and girls in the school, it should be fairly easy to see how many would be without a date

OpenStudy (anonymous):

Using guess and check, multiply 90 and 110 by the same number so that when you add the products up, it equals 16000. Once you find the total amount of boys and girls in the school, it should be fairly easy to see how many would be without a date

OpenStudy (anonymous):

We have a ratio of \(90:110\), but we can simplify that to \(9:11\). The total number of enrollment is \(16,000\). This is a trial-and-error method, and I just list all the factors of 16,000 and multiply it by 9 & 10. And I got \( 16000 \div 20=800\). \[\huge \frac{ 9 \times 800 }{ 11 \times 800 }=\frac{ 7,200 }{ 8,800 }\frac{ men }{ women}\]

OpenStudy (anonymous):

So there are \(7,200\) men and \(8,800\) women. Just subtract the number of men to the number of women. \(\huge 8,800-7,200=1,600\). Therefore, \(1,600\) women would be lonely because they wouldn't have dates for the dance. Poor girls!